A Bibliography of Publications by, and about, Per Brinch Hansen

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

26 July 2022 Version 1.05

Title word cross-reference

N [Bri91c].

-Body [Bri91c].

1971 [FGR71]. 1977 [Whi78].

2000 [Smi01].

3003 [BH66]. 365 [Bri00, Leó00].

4000 [BK66, Bri67a, Bri67b, Bri69a, Bri69b, Bri69c, Bri96s, Bri01c, Bri02i].

5000 [Oli01].

6 [SS01].

701 [Ryc01].

8.50 [Bri77b].

academic [Bri77b]. Access [Fra01]. Achievements [Pin02]. Ada [BBG^{+}81l]. Address [Bri79b, Bri96g]. Akyildiz [Pin02]. Algorithm [Bri95d]. All-Pairs [Bri90a, Bri95a]. Amendment [Ber98]. Amoeba [TRS^{+}01]. Analysis [Bri71a, Bri89a]. Annealing [Bri92g]. Arbitrary [Reg84]. Architecture [Bri77a, Pin02, Whi78]. Architectures [Bri78b, Bri78c]. Arrays [Bri99h]. Atlas [KPH01]. August [BS76, FGR71]. Author [Leó00, Bri79a]. Automata [Bri92d, Bri93d, Bri96p]. Award [Pin02, Pin02].

B [Oli01]. Back [BBG^{+}81a]. Balancing [Bri90b, Bri95b]. Based [Fra01, Lau01]. Batch [Bri01a]. Better [VBM^{+}02]. Biographies [LH03]. BKS [Smi01]. block [Bri94b]. block-structured [Bri94b]. Body [Bri91c]. Book [Bri77b, Kru97, Van95, Whi78]. Books [Nau75]. Bookshelf [VBM^{+}02]. Brinch [Ber98, CHP72, Hor07, Kru97, LH03, Off98, Pin02, Van95, Whi78]. Building [VBM^{+}02]. Bulletproof [VBM^{+}02].

C [Bri00, Leó00]. Call [SP01]. Calls [Bri02f]. Capacity [Lyn01]. Carlo [Bri92e]. Cellular [Bri92d, Bri93d, Bri96p]. Center [Lyn01]. Choice [Bri99e, Bri99i]. Classes [Bri73c, Bri76e, Bri96v, Bri96x, Bri99a, Bri01e, Bri02j, Bri02l]. Classic [Bri01a]. COBOL [BH66]. Collection [Pin02]. Comments [Bri73b, CHP72]. Communicating [Hoa02a]. Comparison [Bri72a, Bri73b, BH79, CHP72]. Compilation [BBG^{+}81k]. Compile [Bri76a]. Compiler [BH66]. Compilers [Bri85a]. complete [Bri69c]. Computational [Bri92d, Bri93b, Bri93d, Bri96h, Bri96p, Van95]. Computer [BK66, Bri82, Pin02, RDH^{+}01, Ryc01, SS01, Bri67a, Bri69a, Pin02]. Computers [Bri85b, Bri83]. Computing [Bri96z, Bri02n, Lyn01, Bri94e]. Concept [Bri78a, Bri96c, Bri02b, Hoa02b]. Concepts [Bri72a, Bri73a, Bri73b, CHP72]. Concerns [Off98]. Concurrency [Bri00, Leó00]. Concurrent [Bri73a, Bri77a, Bri77d, Bri78a, Bri78b, Bri79b, Bri96c, Bri96e, Bri96g, Bri02b, Bri02d, Bri02f, Whi78, Bri78c, Bri75c, Bri75a, Bri75b, Bri75d, Bri76c, Bri93c, Bri96i, Bri96j, Bri96r, Bri96w, Bri01d, Bri02h, Bri02k]. Congress [FGR71]. Conjugate [Bri98a]. Connection [Pin02, BMR01]. Conquer [Bri91d]. Considerations [BK66]. Control [Bri67b, Bri96s, Bri94b]. Cooperating [Dij02a]. correction [Bri79a]. Correspondence [Bri99l]. Course [Bri96o, Bri02g, Bri72b]. courses [Bri83].

D [Bri77b]. Declarations [BBG^{+}81b]. Defense [BBG^{+}81l]. Department [BBG^{+}81l]. Dependent [BBG^{+}81m]. Description [Lyn01]. Design [BK66, Bri74, Bri77c, Bri96a, Bri96b, Bri02a, Lis01, SP01, Bri81a]. Designers [VBM^{+}02]. Digital [Pin02]. Disc [Fra01]. Disc-Based [Fra01]. Disk [Bri76a]. Distributed [Bri78a, Bri87a, Bri96c, Bri96f, Bri01a, Bri02b, Bri02e, SMB01, TRS^{+}01]. Divide [Bri91d]. Division [Bri92b, Bri94c, Bri95d, Reg84]. Do [Bri91a, Bri94a]. Document [BBG^{+}81l]. Dog [Bri76b, Bri96q].

Edison [Bri81a, Bri81b, Bri81c, Bri96a]. edited [Bri77b]. Editor [Bri98d, Bri00]. Editorial [Bri81d]. Efficient [Bri95c, Bri96d, Bri02c, Bri00, Leó00, RDLG99]. Egdon [BHJV01]. Elements [BBG^{+}81h]. Emerging [Dij02d]. Engineering [Pin02]. Engineers [Pin02]. Environment [SMB01]. Equation [Bri92c]. Equations [Bri92a, Bri98a]. Essays [Ber98, Bri96u, Bri97, Kru97]. Evaluation [Bri98b, Bri98c]. Everyone [Bri99f]. Evolution [Bri01b]. Examples [Reg84]. Exceptions [BBG^{+}81c]. Exclusion [BS78]. Expanded [Pin02]. Experience [Bri77d, Bri96e, Bri00, Leó00]. Experiences [TRS^{+}01]. Experimental [CMDD01, SS01]. Expressions [BBG^{+}81i].

F [Bri00, Leó00, Pin02]. Fast [Lyn01, Bri93a]. Faster [Bri91a, Bri94a]. Features [BBG^{+}81m]. File [DN01, Fra01, SMB01]. Files [Bri99m]. Filing [SS01]. Folding [Bri90b]. Fortran [Bri98b, Off98]. Foundations [FGR71]. Fourier [Bri93a]. Front [BBG^{+}81d].

García [Bri00, Leó00]. General [DN01, Ryc01]. General-Purpose [DN01]. Generic [BBG^{+}81e, Bri91b]. Goode [Pin02]. gradient [Bri98a]. Guest [Bri81d].

H [Bri77b]. Hall [Whi78]. Handbook [Bri85b]. Hansen [Ber98, Whi78, CHP72, Hor07, Kru97, LH03, Off98, Pin02, Van95]. held [BS76]. Hennessy [Pin02]. Hierarchical [Dij02b]. hierarchies [BS76]. High [Lyn01, Bri98b, Off98]. History [BG96, Bri96i, Bri93c, Bri96j]. Honors [Pin02]. Householder [Bri90c, Bri92a]. Huxtable [Bri77b]. Hypercubes [Bri91a, Bri94a].

Ian [Pin02]. IBM [Bri85b, Ryc01]. IEEE [Pin02]. IFIP [FGR71]. II [BG96]. Implementation [BBG^{+}81m, BS78, Bri87b, Bri88, Bri89c, Bri96k]. Information [FGR71]. Input [BBG^{+}81f, SS01]. Input-Output [BBG^{+}81f]. Input/Output [SS01]. Insecure [Bri99l, Bri99b]. Integer [Reg84]. Integrity [Fra01]. Interface [Bri76d, SIKH01, Bri98c]. interfaces [BS76]. Interference [Bri94b]. international [BS76]. Introduction [BBG^{+}81g]. Invention [Bri02d]. Issues [BBG^{+}81k].

Java [Bri99b, Bri99f, Bri99l]. Job [Bri76d]. Journals [Nau75]. Joyce [Bri87b, Bri87a, BA88, Bri88, Bri89b, Bri89c, Bri96f, Bri96k, Bri02e]. July [BS76].

KDF9 [BHJV01]. Keynote [Bri79b, Bri96g].

L [Bri00, Leó00]. Laboratories [Ryc01]. Language [BBG^{+}81l, Bri87a, Bri89b, Bri94d, Bri96f, Bri96r, Bri96z, Bri02e, Bri02h, Bri02n, Bri75a, Bri75b, Bri81b, Bri94b, Bri94e, BS76]. languages [BG96]. Laplace [Bri92c]. Large [Lau01]. Leading [Pin02]. Learning [VBM^{+}02]. Length [Bri92b, Bri94c]. León [Bri00]. Letter [Bri98d, Bri00]. Lexical [BBG^{+}81h]. libraries [Bri00, Leó00, RDLG99]. Library [Pin02]. Linear [Bri90d, Bri92a, DF89, Bri98a]. Linger [Bri95d]. Ljubljana [FGR71]. logical [Bri67a]. London [Bri77b]. Long [Bri95d].

M [Bri77b]. Machine [LS01]. Machines [Bri91a, Bri94a]. Manual [BBG^{+}81l, Bri69a]. Marktoberdorf [BS76]. Matrices [Bri99c]. Matter [BBG^{+}81a, BBG^{+}81d]. McKeag [Bri77b]. Mechanism [SP01]. Mergesort [Bri89a]. message [Bri00, Leó00, RDLG99, Bri98c]. Message-Passing [Bri98c]. Methodology [Bri74, Bri93b, Bri96h, Bri00, Leó00, RDLG99]. Methods [Bri99d]. Microcomputer [BH79]. Mills [Bri95d, Pin02]. Minefield [Bri92b, Bri94c]. Model [Bri92d, Bri93b, Bri93d, Bri96h, Bri96p]. Modular [Bri77d, Bri96e]. Monitors [Bri76e, Bri93c, Bri96i, Bri96j, Bri96t, Bri96x, Bri01e, Bri02l, Hoa02b, Bri78e]. Monte [Bri92e]. Motors [Ryc01]. Multi [Fra01]. Multi-Access [Fra01]. Multicomputers [Bri93b, Bri96h]. Multiple [Bri92b, Bri94c, Bri99e]. Multiple-Length [Bri92b, Bri94c]. Multiplication [Bri91b]. Multiprecision [Reg84]. Multiprocessor [Bri78b, Bri78c, Bri78d, BA88, Bri88, Bri89c, Bri96k, Bri96m, Bri81b]. Multiprogramming [Bri69b, Bri69c, Bri70, Bri72d, Bri73d, Bri96n, Bri96y, Bri96-27, Bri01c, Bri02i, Bri02m, Dij01, Dij02c, Bri71b, Bri72c]. Mutual [BS78].

Named [Pin02]. Names [BBG^{+}81i]. Nature [Bri90e, Bri96l, Bri89d]. Network [Bri96m, Bri78d]. Newcastle [BMR01]. News [Van95]. No [VBM^{+}02, Bri77b]. Nook [Kru97]. Nucleus [Bri70, Bri96n]. Numerical [Bri92c].

Obituary [Hor07]. Open [LS01]. Operating [Bri74, Bri76c, Bri76d, Bri76e, BF80, Bri96o, Bri96w, Bri96x, Bri01a, Bri01b, Bri01d, Bri01e, Bri02g, Bri02k, Bri02l, Hoa02b, LS01, Lau01, Lis01, Oli01, RDH^{+}01, SS01, TRS^{+}01, Bri72b, Bri77b, Bri83]. Ordering [Dij02b]. Origin [Bri02f]. Outline [Bri96o, Bri02g, Bri72b]. Output [BBG^{+}81f, SS01]. Overview [SIKH01].

Packages [BBG^{+}81j]. pages [Bri77b]. Pairs [Bri90a, Bri95a]. Paradigms [Van95]. Parallel [Ber98, Bri89a, Bri90e, Bri91d, Bri92d, Bri92e, Bri93d, Bri95c, Bri96d, Bri96l, Bri96p, Bri96z, Bri96u, Bri97, Bri02c, Bri02n, Hoa02c, Kru97, Van95, Bri89d, Bri94b, Bri94e]. Parallelism [Bri99l, RDLG99, Bri99b, Bri00, Leó00]. PASCAL [Bri75c, Bri75a, Bri75b, Bri75d, Bri76c, Bri85a, Bri85b, Bri93c, Bri96i, Bri96j, Bri96r, Bri96w, Bri01d, Bri02h, Bri02k]. passing [Bri00, Leó00, RDLG99, Bri98c]. Past [VBM^{+}02]. People [VBM^{+}02]. Performance [BA88, Bri98b, Off98]. Personal [Bri82, Bri85b, Bri96i, RDH^{+}01, Bri83, Bri93c, Bri96j]. Philco [Smi01]. Philco-2000 [Smi01]. Pilot [RDH^{+}01]. Pioneering [Pin02]. Pipeline [Bri90a, Bri90b, Bri91c, Bri91b, Bri95a, Bri95b]. Practice [Bri00, Leó00, LMW79]. Prentice [Whi78]. Prentice-Hall [Whi78]. President [Pin02]. press [Bri77b]. price [Bri77b]. Primality [Bri92f]. Principles [Bri77c, Bri96b, Bri96o, Bri02a, Bri02g, Bri72b]. Procedure [Bri02f, SP01]. Proceedings [FGR71]. Processes [Bri76e, Bri78a, Bri96c, Bri96x, Bri01e, Bri02b, Bri02l, Dij02a, Dij02b, Hoa02a]. Processing [Bri01a, FGR71]. Program [BBG^{+}81e, BBG^{+}81k, Bri76c, Bri78d, Bri92d, Bri93d, Bri96m, Bri96p, Bri96w, Bri01d, Bri02k]. Programmer [Bri76b, Bri96q, Bri04]. Programming [BBG^{+}81l, Bri73a, Bri74, Bri77d, Bri78a, Bri79b, Bri82, Bri87a, Bri90e, Bri93b, Bri94d, Bri96c, Bri96e, Bri96f, Bri96g, Bri96h, Bri96l, Bri96r, Bri96u, Bri97, Bri99f, Bri02b, Bri02d, Bri02e, Bri02f, Bri02h, Hoa02c, Kru97, LMW79, Van95, BG96, Bri75a, Bri75b, Bri89d, Ber98]. Programs [Bri77a, Bri78b, Bri81c, Bri93b, Bri96h, Bri99j, Bri78c, Whi78]. Proposed [BBG^{+}81l]. Publication [Bri96z, Bri02n, Bri94e]. Pulawy [Bri67b, Bri96s]. Purpose [DN01, Bri75c].

R [Bri77b]. Radix [Reg84]. Ratio [Bri71a]. RC [BK66, Bri67a, Bri67b, Bri69a, Bri69b, Bri69c, Bri96s, Bri01c, Bri02i]. RC-4000 [Bri69a]. Real [Bri67b, Bri96s]. Real-Time [Bri67b, Bri96s]. Receives [Pin02]. Recognized [Pin02]. Recognizes [Pin02]. Recursion [Bri95c, Bri96d, Bri99g, Bri02c, Bri00, Leó00, RDLG99]. Rediscovered [Bri90d]. Reduction [Bri92a, Bri90c]. Reference [BBG^{+}81l, Bri69a]. Reliable [SP01]. Remote [Bri02f, SP01]. Repetition [Bri99k, Bri99n]. Reply [Bri73b]. Report [Bri89b]. Representation [BBG^{+}81m]. Reproducible [Bri78e, Bri96t]. Research [Ryc01]. Response [Bri71a, Leó00]. Review [Ber98, Bri77b, Bri97, Whi78]. Reviews [Van95, Whi78]. Revisited [Bri92b, Bri94c, DF89]. Rodríguez [Bri00]. Rules [BBG^{+}81q].

Sande [Bri00, Leó00]. Scheduling [Bri71a, Bri76a, Bri71b, Bri72c]. school [BS76]. Science [Bri92d, Bri93b, Bri93d, Bri96h, Bri96p, Van95]. Scientific [Bri96z, Bri02n, Bri94e]. Search [Ber98, Bri90d, Bri96u, Bri97, DF89, Kru97].

Secondary [DN01]. Semaphore [Lau01]. Semaphores [Bri02f]. Sequential [Dij02a, Dij02b, Hoa02a]. Shared [Bri73c, Bri96v, Bri02j, SMB01]. Sharing [CMDD01, RT01]. Short [Bri71b, Bri72c]. Short-term [Bri71b, Bri72c]. Siemens [BH66]. Simple [Bri99h, Bri99i, Bri99j, Bri99k, SMB01]. Simplicity [Ber98, Bri96u, Bri97, Kru97]. Simulated [Bri92g]. Single [LS01]. Single-User [LS01]. Small [SS01]. Society [Pin02]. Software [Bri69b, Bri69c, Bri95e, Bri95f, Bri01c, Bri02i, Pin02, VBM^{+}02]. Solo [Bri76c, Bri76d, Bri76e, Bri96w, Bri96x, Bri01d, Bri01e, Bri02k, Bri02l]. Solution [Bri92c, Bri98a]. Sort [Bri91a, Bri94a]. Specification [BS78]. Specifications [BBG^{+}81m]. Standard [BBG^{+}81l]. Stanford [Pin02]. Star [SIKH01]. Statements [BBG^{+}81n]. States [BBG^{+}81l]. Storage [DN01]. Story [Bri04]. Strategies [VBM^{+}02]. Structure [BBG^{+}81k, Dij01, Dij02c, Bri67a]. Structured [Bri72d, Bri96y, Bri02m, LMW79, Bri94b]. Structuring [Hoa02b]. Studies [Van95, Bri77b]. Subprograms [BBG^{+}81o]. summer [BS76]. SuperPascal [Bri94b, Bri94d, Bri94e, Bri95e, Bri95f, Bri96z, Bri02n]. Supervisor [KPH01]. Synchronizing [Bri72a, Bri73b, CHP72]. Synthesis [Dij02d]. System [Bri67b, Bri69b, Bri69c, Bri70, Bri73d, Bri74, Bri76c, Bri76d, Bri76e, BF80, Bri96n, Bri96o, Bri96s, Bri96w, Bri96x, Bri96-27, Bri01c, Bri01d, Bri01e, Bri02g, Bri02i, Bri02k, Bri02l, BHJV01, CMDD01, DN01, Dij01, Dij02c, Fra01, Hoa02b, LS01, Lau01, Lis01, Oli01, RDH^{+}01, RT01, Smi01, SS01, SMB01, TRS^{+}01, VBM^{+}02, Bri72b, Bri83]. Systems [Bri87a, Bri96f, Bri01a, Bri01b, Bri02e, FGR71, Bri71b, Bri72c, Bri77b].

Tasks [BBG^{+}81p]. Technical [Bri99l]. term [Bri71b, Bri72c]. Testing [Bri73d, Bri92f, Bri96t, Bri96-27, Bri78e]. Text [Bri99m]. Theory [Hoa02c, LMW79]. Time [Bri67b, Bri76a, CMDD01, RT01, Bri96s]. Time-Sharing [CMDD01, RT01]. Tour [Bri92b, Bri94c]. Transform [Bri93a]. Treasure [VBM^{+}02]. Tree [Bri91a, Bri94a]. Trials [Bri92e]. Trio [BF80]. Turnaround [Lyn01]. Two [Bri72a, Bri73b, CHP72]. Types [BBG^{+}81b, Bri75d].

Unbounded [Bri99n]. Unite [BMR01]. United [BBG^{+}81l]. Units [BBG^{+}81e]. Universal [Bri75d]. University [Lyn01]. Unix [RT01]. Unixes [BMR01]. User [LS01, SIKH01]. Using [Bri83, Reg84].

Venus [Lis01]. Vest [VBM^{+}02]. Visibility [BBG^{+}81q]. Volume [FGR71].

WFS [SMB01]. Wilson [Bri77b]. Winner [Pin02]. Witt [Bri95d]. Works [Pin02]. World [BMR01].

Young [Bri76b, Bri96q]. Yugoslavia [FGR71].

References

BrinchHansen:1988:JPM

[BA88] Per Brinch Hansen and Rangachari Anand. Joyce performance on a multiprocessor. Electrical Engineering and Computer Science Technical Report 28, School of Computer and Information Science, Syracuse University, Syracuse, NY, USA, September 1988. ii + 19

pp. URL https://surface.syr.edu/eecs_techreports/28.

Brauer:1981:BM

[BBG^{+}81a] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Back matter. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 207–243. ISBN 3-540-10693-6. URL https://

link.springer.com/content/pdf/bbm:978-3-540-38620-9/1.

Brauer:1981:DT

[BBG^{+}81b] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Declarations and types. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 15–45. ISBN 3-540-10693-6. URL https:

//link.springer.com/chapter/10.1007/3-540-10693-6_3.

Brauer:1981:E

[BBG^{+}81c] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Exceptions. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 143–153. ISBN 3-540-10693-6. URL https://

link.springer.com/chapter/10.1007/3-540-10693-6_11.

Brauer:1981:FM

[BBG^{+}81d] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Front matter. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages i–x. ISBN 3-540-10693-6. URL https://link.

springer.com/content/pdf/bfm:978-3-540-38620-9/1.

Brauer:1981:GPU

[BBG^{+}81e] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Generic program units. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 155–167. ISBN 3-540-10693-6. URL https://

link.springer.com/chapter/10.1007/3-540-10693-6_12.

Brauer:1981:IO

[BBG^{+}81f] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Input-output. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 183–206. ISBN 3-540-10693-6. URL https://

link.springer.com/chapter/10.1007/3-540-10693-6_14.

Brauer:1981:I

[BBG^{+}81g] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Introduction. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 1–6. ISBN 3-540-10693-6. URL https://link.

springer.com/chapter/10.1007/3-540-10693-6_1.

Brauer:1981:LE

[BBG^{+}81h] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Lexical elements. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 7–13. ISBN 3-540-10693-6. URL https:/

/link.springer.com/chapter/10.1007/3-540-10693-6_2.

Brauer:1981:NE

[BBG^{+}81i] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Names and expressions. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 47–71. ISBN 3-540-10693-6. URL https:

//link.springer.com/chapter/10.1007/3-540-10693-6_4.

Brauer:1981:P

[BBG^{+}81j] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Packages. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 93–103. ISBN 3-540-10693-6. URL https://link.springer.com/

chapter/10.1007/3-540-10693-6_7.

Brauer:1981:PSC

[BBG^{+}81k] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Program structure and compilation issues. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 131–141. ISBN 3-540-

10693-6. URL https://link.springer.com/chapter/10.1007/
3-540-10693-6_10.

Brauer:1981:PLA

[BBG^{+}81l] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth, editors. The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense, volume 106 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. ISBN 3-540-10693-6.

Brauer:1981:RSI

[BBG^{+}81m] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Representation specifications and implementation dependent features. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 169–181. ISBN 3-540-10693-6. URL https://link.springer.

com/chapter/10.1007/3-540-10693-6_13.

Brauer:1981:Sa

[BBG^{+}81n] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Statements. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 73–82. ISBN 3-540-10693-6. URL https:

//link.springer.com/chapter/10.1007/3-540-10693-6_5.

Brauer:1981:Sb

[BBG^{+}81o] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Subprograms. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 83–92. ISBN 3-540-10693-6. URL https:

//link.springer.com/chapter/10.1007/3-540-10693-6_6.

Brauer:1981:T

[BBG^{+}81p] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Tasks. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 115–130. ISBN 3-540-10693-6. URL https://link.springer.

com/chapter/10.1007/3-540-10693-6_9.

Brauer:1981:VR

[BBG^{+}81q] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth. Visibility rules. In The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense [BBG^{+}81l], pages 105–114. ISBN 3-540-10693-6. URL https://

link.springer.com/chapter/10.1007/3-540-10693-6_8.

Berman:1998:ARP

[Ber98] A. Michael Berman. Amendment to review of Per Brinch Hansen, The Search for Simplicity — Essays in Parallel Programming. ACM SIGPLAN Notices, 33(3):36, March 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

BrinchHansen:1980:TOS

[BF80] Per Brinch Hansen and Jon Fellows. The Trio operating system. Software—Practice and Experience, 10(11):943–948, November 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bergin:1996:HPL

[BG96] T. J. Bergin, Jr. and R. G. Gibson, Jr., editors. History of programming languages: II. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-201-89502-1. xvi + 864 pp. LCCN QA76.7 .H57

BrinchHansen:1966:CCS

[BH66] Per Brinch Hansen and Roger House. The COBOL compiler for the Siemens 3003. Nordisk tidskrift for informationsbehandling, 6(1):1–23, March 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://brinch-hansen.net/papers/1966a.pdf;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=
0006-3835&volume=6&issue=1&spage=1.

BrinchHansen:1979:MC

[BH79] Per Brinch Hansen and Charles Hayden. Microcomputer comparison. Software—Practice and Experience, 9(3):211–217, March 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Burns:2001:ESK

[BHJV01] David Burns, E. Neville Hawkins, D. Robin Judd, and John L. Venn. The Egdon system for the KDF9. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 102–114. ISBN 0-387-95113-X. LCCN

QA76.76.O63 B7425 2001. URL http://www.loc.gov/catdir/
enhancements/fy0816/00045036-d.html;   http://www.loc.
gov/catdir/enhancements/fy0816/00045036-t.html;   https:
//link.springer.com/book/10.1007/978-1-4757-3510-9.

BrinchHansen:1966:DCR

[BK66] Per Brinch Hansen and Peter Kraft. Design considerations for the RC 4000 computer. Report, A/S Regnecentralen, Copenhagen, Denmark, March 1966. 46 pp. URL http://brinch-hansen.

net/papers/1966b.pdf.

Brownbridge:2001:NCU

[BMR01] David R. Brownbridge, Lindsay F. Marshall, and Brian Randell. The Newcastle Connection or Unixes of the world unite. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 528–549. ISBN 0-387-95113-X. LCCN

QA76.76.O63 B7425 2001. URL http://www.loc.gov/catdir/
enhancements/fy0816/00045036-d.html;   http://www.loc.
gov/catdir/enhancements/fy0816/00045036-t.html;   https:
//link.springer.com/book/10.1007/978-1-4757-3510-9.

BrinchHansen:1967:LSR

[Bri67a] Per Brinch Hansen. The logical structure of the RC 4000 computer. BIT, 7(3):191–199, September 1967. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://brinch-hansen.net/papers/1967a.pdf;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=
0006-3835&volume=7&issue=3&spage=191.

BrinchHansen:1967:RRT

[Bri67b] Per Brinch Hansen. The RC 4000 real-time control system at Pulawy. BIT, 7(4):279–288, December 1967. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://brinch-hansen.net/papers/1967b.pdf;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=
0006-3835&volume=7&issue=4&spage=279.

BrinchHansen:1969:RCR

[Bri69a]  Per Brinch Hansen.  RC-4000 Computer Reference Man-

ual. A/S Regnecentralen, Copenhagen, Denmark, second edition, June 1969. 85 pp. URL http://bitsavers.org/

pdf/regnecentralen/RC_4000_Reference_Manual_Jun69.pdf;
 http://brinch-hansen.net/papers/1969b.pdf;   http://www.
math.utah.edu/~beebe/RC-4000. RCSL No: 55-D1.

BrinchHansen:1969:RSMa

[Bri69b] Per Brinch Hansen. RC 4000 software: Multiprogramming system. In RC 4000 Software: Multiprogramming System, Part I: General Description, pages 13–52. Regnecentralen, Copenhagen, Denmark,

1969. URL http://brinch-hansen.net/papers/1969a.pdf.

BrinchHansen:1969:RSMb

[Bri69c] Per Brinch Hansen. RC 4000 Software Multiprogramming System (complete). A/S Regnecentralen, Copenhagen, Denmark, second edition, April 1969. 159 pp. URL http:

//brinch-hansen.net/papers/1969b.pdf;   http://brinch-
hansen.net/papers/1969c.pdf. RCSL No: 55-D17.

BrinchHansen:1970:NMS

[Bri70] Per Brinch Hansen. The nucleus of a multiprogramming system. Communications of the Association for Computing Machinery, 13(4):238–241, 250, April 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://brinch-

hansen.net/papers/1970a.pdf.

BrinchHansen:1971:ARR

[Bri71a] Per Brinch Hansen. An analysis of response ratio scheduling. In Freiman et al. [FGR71], pages 479–484. ISBN 0-7204-2063-6 (two volumes). LCCN QA75.5 .I532 1971, TK7885.A1 I61 1971.

BrinchHansen:1971:STS

[Bri71b] Per Brinch Hansen. Short-term scheduling in multiprogramming systems. In Proceedings of the third symposium on Operating systems principles — SOSP ’71. ACM Press, New York, NY 10036, USA, 1971.

BrinchHansen:1972:CTS

[Bri72a] Per Brinch Hansen. A comparison of two synchronizing concepts. Acta Informatica, 1(3):190–199, February 1972. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See comments [CHP72] and reply [Bri73b].

BrinchHansen:1972:OCO

[Bri72b] Per Brinch Hansen. An outline of a course on operating system principles. In C. A. R. Hoare and R. H. Perrott, editors, Operating Systems Techniques, Proceedings of a Seminar at Queen’s University, Belfast, Northern Ireland, August–September 1971, pages 29–36. Academic Press, New York, USA, 1972. URL http://

brinch-hansen.net/papers/1971a.pdf.

BrinchHansen:1972:STS

[Bri72c] Per Brinch Hansen. Short-term scheduling in multiprogramming systems. Operating Systems Review, 6(1–2):101–105, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

BrinchHansen:1972:SM

[Bri72d] Per Brinch Hansen. Structured multiprogramming. Communications of the Association for Computing Machinery, 15(7):574–578, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317

(electronic). URL http://brinch-hansen.net/papers/1972a.
pdf.

BrinchHansen:1973:CPC

[Bri73a] Per Brinch Hansen. Concurrent programming concepts. ACM Computing Surveys, 5(4):223–245, December 1973. CODEN CMSVAN. ISSN 0010-4892.

BrinchHansen:1973:RCC

[Bri73b] Per Brinch Hansen. A reply to comments on A Comparison of Two Synchronizing Concepts. Acta Informatica, 2(??):189–190, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [Bri72a, CHP72].

BrinchHansen:1973:SC

[Bri73c] Per Brinch Hansen. Shared classes. In Operating System Principles, pages 226–232. Prentice-Hall, Upper Saddle River, NJ 07458,

USA, 1973. URL http://brinch-hansen.net/papers/1973b.
pdf.

BrinchHansen:1973:TMS

[Bri73d] Per Brinch Hansen. Testing a multiprogramming system. Software —Practice and Experience, 3(2):145–150, April/June 1973. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

URL http://brinch-hansen.net/papers/1973c.pdf.

BrinchHansen:1974:PMO

[Bri74] Per Brinch Hansen. A programming methodology for operating system design. In Jack L. Rosenfeld, editor, Information Processing, Proceedings of the 6th IFIP Congress 1974, Stockholm, Sweden, August 5–10, 1974, pages 394–397. North-Holland Publishing Co., Amsterdam, The Netherlands, 1974.

BrinchHansen:1975:PLCa

[Bri75a] Per Brinch Hansen. The programming language Concurrent Pascal. IEEE Transactions on Software Engineering, SE-1(2):199–207, June 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://brinch-

hansen.net/papers/1975a.pdf;   http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=6312840.

BrinchHansen:1975:PLCb

[Bri75b] Per Brinch Hansen. The programming language Concurrent Pascal. In Friedrich L. Bauer and Klaus Samelson, editors, Language Hierarchies and Interfaces, International Summer School, Marktoberdorf, Germany, July 23–August 2, 1975, volume 46 of Lecture Notes in Computer Science, pages 82–110. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975.

BrinchHansen:1975:PCP

[Bri75c] Per Brinch Hansen. The purpose of Concurrent PASCAL. ACM SIGPLAN Notices, 10(6):305–309, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

BrinchHansen:1975:UTC

[Bri75d] Per Brinch Hansen. Universal types in Concurrent Pascal. Information Processing Letters, 3(6):165–166, July ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

BrinchHansen:1976:DSC

[Bri76a] Per Brinch Hansen. Disk scheduling at compile time. Software— Practice and Experience, 6(2):201–205, April/June 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

BrinchHansen:1976:PYD

[Bri76b] Per Brinch Hansen. The programmer as a young dog. In P. Sveistrup, P. Naur, H. B. Hansen, and C. Gram, editors, Niels Ivar Bech — en epoke i edb-udviklingen i Danmark. (Danish) [Niels Ivar Bech — an epoch in data processing development in Denmark], pages 65–68. Data, Copenhagen, Denmark, 1976. URL http://

brinch-hansen.net/papers/1976e.pdf. translation to English

by the author.

BrinchHansen:1976:SOSa

[Bri76c] Per Brinch Hansen. The Solo operating system: a Concurrent Pascal program. Software—Practice and Experience, 6(2): 141–149, April/June 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://brinch-hansen.

net/papers/1976b.pdf.

BrinchHansen:1976:SOSb

[Bri76d] Per Brinch Hansen. The Solo operating system: Job interface. Software—Practice and Experience, 6(2):151–164, April/June 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

BrinchHansen:1976:SOSc

[Bri76e] Per Brinch Hansen. The Solo operating system: Processes, monitors, and classes. Software—Practice and Experience, 6(2): 165–200, April/June 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://brinch-hansen.

net/papers/1976c.pdf.

BrinchHansen:1977:ACP

[Bri77a] Per Brinch Hansen. The Architecture of Concurrent Programs. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1977. ISBN 0-13-044628-9. xvii + 317 pp. LCCN QA76.73.C65B73.

BrinchHansen:1977:BRS

[Bri77b] Per Brinch Hansen. Book review: Studies in operating systems, R. M. McKeag and R. Wilson, edited by D. H. R. Huxtable, academic press, London, 1976. no. of pages: 263. price: £8.50. Software —Practice and Experience, 7(2):302–303, March 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

BrinchHansen:1977:DP

[Bri77c] Per Brinch Hansen. Design principles. In The Architecture of Concurrent Programs [Bri77a], pages 3–14. ISBN 0-13-044628-

9. LCCN QA76.73.C65B73. URL http://brinch-hansen.net/
papers/1977c.pdf.

BrinchHansen:1977:EMC

[Bri77d] Per Brinch Hansen. Experience with modular concurrent programming. IEEE Transactions on Software Engineering, SE-3 (2):156–159, March 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://brinch-hansen.

net/papers/1977a.pdf.

BrinchHansen:1978:DPC

[Bri78a] Per Brinch Hansen. Distributed processes: a concurrent programming concept. Communications of the Association for Computing Machinery, 21(11):934–941, November 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

brinch-hansen.net/papers/1978a.pdf.

BrinchHansen:1978:MACa

[Bri78b] Per Brinch Hansen. Multiprocessor architectures for concurrent programs. In Richard H. Austing, Dennis M. Conti, and Gerald L. Engel, editors, Proceedings of the 1978 ACM Annual Conference, Washington, DC, USA, December 4–6, 1978, Volume I, pages 317– 323. ACM Press, New York, NY 10036, USA, 1978. URL https:

//doi.org/10.1145/800127.804119.

BrinchHansen:1978:MACb

[Bri78c] Per Brinch Hansen. Multiprocessor architectures for concurrent programs. ACM SIGARCH Computer Architecture News, 7(4):4– 23, December 1978. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

BrinchHansen:1978:NMP

[Bri78d] Per Brinch Hansen. Network: a multiprocessor program. IEEE Transactions on Software Engineering, SE-4(3):194–199, May 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://brinch-hansen.net/papers/1977d.pdf.

BrinchHansen:1978:RTM

[Bri78e] Per Brinch Hansen. Reproducible testing of monitors. Software— Practice and Experience, 8(6):721–729, November/December 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://brinch-hansen.net/papers/1978b.pdf.

BrinchHansen:1979:AC

[Bri79a] Per Brinch Hansen. Author’s correction. Software—Practice and Experience, 9(10):879–880, October 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

BrinchHansen:1979:KAC

[Bri79b] Per Brinch Hansen. A keynote address on concurrent programming. Computer, 12(5):50–56, May 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://brinch-

hansen.net/papers/1978c.pdf.

BrinchHansen:1981:DE

[Bri81a] Per Brinch Hansen. The design of Edison. Software—Practice and Experience, 11(4):363–396, April 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://brinch-

hansen.net/papers/1981b.pdf.

BrinchHansen:1981:EML

[Bri81b] Per Brinch Hansen. Edison—a multiprocessor language. Software —Practice and Experience, 11(4):325–361, April 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

BrinchHansen:1981:EP

[Bri81c] Per Brinch Hansen. Edison programs. Software—Practice and Experience, 11(4):397–414, April 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

BrinchHansen:1981:GEG

[Bri81d] Per Brinch Hansen. Guest editorial: Guest editorial. Software— Practice and Experience, 11(4):323, April 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

BrinchHansen:1982:PPC

[Bri82] Per Brinch Hansen. Programming a Personal Computer. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1982. ISBN 0-13-730283-5 (hardcover), 0-13-730283-5 (paperback). xi + 388 pp. LCCN QA76.73.E34 B74 1982.

BrinchHansen:1983:UPC

[Bri83] Per Brinch Hansen. Using personal computers in operating system courses. Operating Systems Review, 17(3):41–44, July 1983. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

BrinchHansen:1985:BHP

[Bri85a] Per Brinch Hansen. Brinch Hansen on Pascal Compilers. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1985. ISBN 0-13-083098-4, 0-13-083122-0 (paperback). x + 310 pp. LCCN QA76.73.P2 B75 1985. US$25.00.

BrinchHansen:1985:PHI

[Bri85b] Per Brinch Hansen. The Pascal Handbook for IBM Personal Computers. ????, ????, 1985. ISBN ???? ???? pp.

BrinchHansen:1987:JPL

[Bri87a] Per Brinch Hansen. Joyce — a programming language for distributed systems. Software—Practice and Experience, 17(1):29–50, January 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-

024X (electronic). URL http://brinch-hansen.net/papers/
1987a.pdf.

BrinchHansen:1987:JI

[Bri87b] Per Brinch Hansen. A Joyce implementation. Software—Practice and Experience, 17(4):267–276, April 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

BrinchHansen:1988:MIJ

[Bri88] Per Brinch Hansen. A multiprocessor implementation of Joyce. Electrical Engineering and Computer Science Technical Report 29, School of Computer and Information Science, Syracuse University, Syracuse, NY, USA, September 1988. ii + 19 pp. URL https://

surface.syr.edu/eecs_techreports/29.

BrinchHansen:1989:APM

[Bri89a] Per Brinch Hansen. Analysis of a parallel mergesort. Electrical Engineering and Computer Science Technical Report 55 (SU-CIS-89-03), School of Computer and Information Science, Syracuse University, Syracuse, NY, USA, July 1989. ii + 17 pp. URL

https://surface.syr.edu/eecs_techreports/55.

BrinchHansen:1989:JLR

[Bri89b] Per Brinch Hansen. The Joyce language report. Software—Practice and Experience, 19(6):553–578, June 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

brinch-hansen.net/papers/1989c.pdf.

BrinchHansen:1989:MIJ

[Bri89c] Per Brinch Hansen. A multiprocessor implementation of Joyce. Software—Practice and Experience, 19(6):579–592, June 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://brinch-hansen.net/papers/1989a.pdf.

BrinchHansen:1989:NPP

[Bri89d] Per Brinch Hansen. The nature of parallel programming. Electrical Engineering and Computer Science Technical Report 59 (SU-CIS-89-08), School of Computer and Information Science, Syracuse University, Syracuse, NY, USA, March 1989. iii + 19 pp. URL

htps://surface.syr.edu/eecs_techreports/59.

BrinchHansen:1990:APP

[Bri90a] Per Brinch Hansen. The all-pairs pipeline. Electrical Engineering and Computer Science Technical Report 77 (SU-CIS-90-40), School of Computer and Information Science, Syracuse University, Syracuse, NY, USA, December 1990. ii + 17 pp. URL https://

surface.syr.edu/eecs_techreports/77.

BrinchHansen:1990:BPF

[Bri90b] Per Brinch Hansen. Balancing a pipeline by folding. Electrical Engineering and Computer Science Technical Report 76 (SU-CIS-90-41), School of Computer and Information Science, Syracuse University, Syracuse, NY, USA, December 1990. iii + 10 pp. URL

https://surface.syr.edu/eecs_techreports/76/.

BrinchHansen:1990:HR

[Bri90c] Per Brinch Hansen. Householder reduction. Electrical Engineering and Computer Science Technical Report 78 (SU-CIS-90-39), School of Computer and Information Science, Syracuse University, Syracuse, NY, USA, December 1990. ii + 13 pp. URL htps://

surface.syr.edu/eecs_techreports/78.

BrinchHansen:1990:LSR

[Bri90d] Per Brinch Hansen. The linear search rediscovered. Structured Programming, 11:53–55, 1990. CODEN STPGEM. ISSN 0935-

1183. URL http://brinch-hansen.net/papers/1990a.pdf. See

[DF89].

BrinchHansen:1990:NPP

[Bri90e] Per Brinch Hansen. The nature of parallel programming. In M. A. Arbib and J. A. Robinson, editors, Natural and Artificial Parallel Computation, pages 31–46. MIT Press, Cambridge, MA, USA,

1990. URL http://brinch-hansen.net/papers/1989b.pdf. In-

augural Symposium at Syracuse University, Syracuse, NY, March 1989.

BrinchHansen:1991:DHS

[Bri91a] Per Brinch Hansen. Do hypercubes sort faster than tree machines? Electrical Engineering and Computer Science Technical Report 132 (SU-CIS-91-44), School of Computer and Information Science, Syracuse University, Syracuse, NY, USA, December 1991. ii + 10

pp. URL https://surface.syr.edu/eecs_techreports/132.

BrinchHansen:1991:GMP

[Bri91b] Per Brinch Hansen. A generic multiplication pipeline. Electrical Engineering and Computer Science Technical Report 111 (SU-CIS-91-21), School of Computer and Information Science, Syracuse University, Syracuse, NY, USA, July 1991. ii + 15 pp. URL

https://surface.syr.edu/eecs_techreports/111.

BrinchHansen:1991:BP

[Bri91c] Per Brinch Hansen. The N-body pipeline. Electrical Engineering and Computer Science Technical Report 120 (SU-CIS-91-09), School of Computer and Information Science, Syracuse University, Syracuse, NY, USA, March 1991. ii + 11 pp. URL https://

surface.syr.edu/eecs_techreports/120.

BrinchHansen:1991:PDC

[Bri91d] Per Brinch Hansen. Parallel divide and conquer. Electrical Engineering and Computer Science Technical Report 131 (SU-CIS-91-45), School of Computer and Information Science, Syracuse University, Syracuse, NY, USA, December 1991. ii + 17 pp. URL

https://surface.syr.edu/eecs_techreports/131.

BrinchHansen:1992:HRL

[Bri92a] Per Brinch Hansen. Householder reduction of linear equations. ACM Computing Surveys, 24(2):185–194, June 1992. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

URL http://brinch-hansen.net/papers/1992a.pdf;   http:
//www.acm.org/pubs/toc/Abstracts/0360-0300/130851.html.

BrinchHansen:1992:MLD

[Bri92b] Per Brinch Hansen. Multiple-length division revisited: a tour of the minefield. Electrical Engineering and Computer Science Technical Report 166 (SU-CIS-92-19), School of Computer and Information Science, Syracuse University, Syracuse, NY, USA, September 1992.

ii + 24 pp. URL https://surface.syr.edu/eecs_techreports/
166.

BrinchHansen:1992:NSL

[Bri92c] Per Brinch Hansen. Numerical solution of Laplace’s equation. Electrical Engineering and Computer Science Technical Report 168 (SU-CIS-92-17), School of Computer and Information Science, Syracuse University, Syracuse, NY, USA, September 1992. ii + 14

pp. URL https://surface.syr.edu/eecs_techreports/168.

BrinchHansen:1992:PCA

[Bri92d] Per Brinch Hansen. Parallel cellular automata: a model program for computational science. Electrical Engineering and Computer Science Technical Report 167 (SU-CIS-92-18), School of Computer and Information Science, Syracuse University, Syracuse, NY, USA, September 1992. ii + 29 pp. URL https://surface.syr.edu/

eecs_techreports/167.

BrinchHansen:1992:PMC

[Bri92e] Per Brinch Hansen. Parallel Monte Carlo trials. Electrical Engineering and Computer Science Technical Report 171 (SU-CIS-92-11), School of Computer and Information Science, Syracuse University, Syracuse, NY, USA, July 1992. ii + 8 pp. URL https://

surface.syr.edu/eecs_techreports/171.

BrinchHansen:1992:PT

[Bri92f] Per Brinch Hansen. Primality testing. Electrical Engineering and Computer Science Technical Report 169 (SU-CIS-92-13), School of Computer and Information Science, Syracuse University, Syracuse, NY, USA, July 1992. ii + 9 pp. URL https://surface.syr.edu/

eecs_techreports/169.

BrinchHansen:1992:SA

[Bri92g] Per Brinch Hansen. Simulated annealing. Electrical Engineering and Computer Science Technical Report 170 (SU-CIS-02-12), School of Computer and Information Science, Syracuse University, Syracuse, NY, USA, July 1992. ??–?? pp. URL https://surface.

syr.edu/eecs_techreports/170.

BrinchHansen:1993:FFT

[Bri93a] Per Brinch Hansen. The Fast Fourier Transform. Electrical Engineering and Computer Science Technical Report 130 (SU-CIS-91-46), School of Computer and Information Science, Syracuse University, Syracuse, NY, USA, December 1993. ii + 21 pp. URL

https://surface.syr.edu/eecs_techreports/130.

BrinchHansen:1993:MPC

[Bri93b] Per Brinch Hansen. Model programs for computational science: a programming methodology for multicomputers. Concurrency: Practice and Experience, 5(5):407–423, August 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL

http://brinch-hansen.net/papers/1993b.pdf.

BrinchHansen:1993:MCP

[Bri93c] Per Brinch Hansen. Monitors and Concurrent Pascal: a personal history. ACM SIGPLAN Notices, 28(3):1–35, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://brinch-

hansen.net/papers/1993a.pdfa;   http://www.acm.org:80/
pubs/citations/proceedings/plan/154766/p1-hansen/.

BrinchHansen:1993:PCA

[Bri93d] Per Brinch Hansen. Parallel cellular automata: a model program for computational science. Concurrency: Practice and Experience, 5(5):425–448, August 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http://brinch-

hansen.net/papers/1993c.pdf.

BrinchHansen:1994:DHS

[Bri94a] Per Brinch Hansen. Do hypercubes sort faster than tree machines? Concurrency: Practice and Experience, 6(2):143–151, April 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

BrinchHansen:1994:ICS

[Bri94b] Per Brinch Hansen. Interference control in SuperPascal — a blockstructured parallel language. The Computer Journal, 37(5):399– 406, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://brinch-hansen.net/papers/
1994d.pdf;   http://comjnl.oxfordjournals.org/content/
37/5/399.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_37/Issue_05/Vol37_05.body.html#AbstractHansen.

BrinchHansen:1994:MLD

[Bri94c] Per Brinch Hansen. Multiple-length division revisited: a tour of the minefield. Software—Practice and Experience, 24(6):579–601, June 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X

(electronic). URL http://brinch-hansen.net/papers/1994b.

pdf. This paper derives an algorithm for division of long integers, and implements it as a literate program, although without identifier cross-references. See also related work on division [Bri95d, Reg84] .

BrinchHansen:1994:PLS

[Bri94d] Per Brinch Hansen. The programming language SuperPascal. Software—Practice and Experience, 24(5):467–483, May 1994. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

BrinchHansen:1994:SPL

[Bri94e] Per Brinch Hansen. SuperPascal — a publication language for parallel scientific computing. Concurrency: Practice and Experience, 6(5):461–483, August 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http://brinch-

hansen.net/papers/1994c.pdf.

BrinchHansen:1995:APP

[Bri95a] Per Brinch Hansen. The all-pairs pipeline. In Studies in Computational Science, chapter 4, pages 73–89. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. URL http://brinch-hansen.net/

papers/1995d.pdf.

BrinchHansen:1995:BP

[Bri95b] Per Brinch Hansen. Balancing a pipeline. In Studies in Computational Science, chapter 5, pages 90–100. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. URL http://brinch-hansen.

net/papers/1995e.pdf.

BrinchHansen:1995:EPR

[Bri95c] Per Brinch Hansen. Efficient parallel recursion. ACM SIGPLAN Notices, 30(12):9–16, December 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

http://brinch-hansen.net/papers/1995c.pdf.

BrinchHansen:1995:LDA

[Bri95d] Per Brinch Hansen. The long division algorithm of Linger, Mills and Witt. Software—Practice and Experience, 25(1):109, January 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See [LMW79, Bri94c].

BrinchHansen:1995:SSa

[Bri95e] Per Brinch Hansen. The SuperPascal software. ACM SIGNUM Newsletter, 30(3):27, July 1995. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

BrinchHansen:1995:SSb

[Bri95f] Per Brinch Hansen. The SuperPascal software. ACM SIGSOFT Software Engineering Notes, 20(5):91, December 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

BrinchHansen:1996:DE

[Bri96a] Per Brinch Hansen. Design of Edison. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 17, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:DP

[Bri96b] Per Brinch Hansen. Design principles. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 12, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:DPC

[Bri96c] Per Brinch Hansen. Distributed processes: a concurrent programming concept. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 14, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:EPR

[Bri96d] Per Brinch Hansen. Efficient parallel recursion. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 25, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:EMC

[Bri96e] Per Brinch Hansen. Experience with modular concurrent programming. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 11, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:JPL

[Bri96f] Per Brinch Hansen. Joyce; a programming language for distributed systems. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 18, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:KAC

[Bri96g] Per Brinch Hansen. Keynote address on concurrent programming. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 16, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:MPC

[Bri96h] Per Brinch Hansen. Model programs for computational science: a programming methodology for multicomputers. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 22, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:MCPa

[Bri96i] Per Brinch Hansen. Monitors and Concurrent Pascal: a personal history. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 21, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:MCPb

[Bri96j] Per Brinch Hansen. Monitors and Concurrent Pascal: a personal history. In Bergin, Jr. and Gibson, Jr. [BG96], pages 121–172. ISBN 0-201-89502-1. LCCN QA76.7 .H57 1996.

BrinchHansen:1996:MIJ

[Bri96k] Per Brinch Hansen. Multiprocessor implementation of Joyce. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 19, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:NPP

[Bri96l] Per Brinch Hansen. Nature of parallel programming. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 20, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:NMP

[Bri96m] Per Brinch Hansen. Network: a multiprocessor program. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 13, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:NMS

[Bri96n] Per Brinch Hansen. Nucleus of a multiprogramming system. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 2, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:OCO

[Bri96o] Per Brinch Hansen. Outline of a course on operating system principles. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 3, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:PCA

[Bri96p] Per Brinch Hansen. Parallel cellular automata: a model program for computational science. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 23, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:PYD

[Bri96q] Per Brinch Hansen. Programmer as a young dog. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 10, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:PLC

[Bri96r] Per Brinch Hansen. Programming language Concurrent Pascal. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 7, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:RRT

[Bri96s] Per Brinch Hansen. RC 4000 real-time control system at Pulawy. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 1, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:RTM

[Bri96t] Per Brinch Hansen. Reproducible testing of monitors. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 15, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:SSE

[Bri96u] Per Brinch Hansen. The Search for Simplicity: Essays in Parallel Programming. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7566-7. xiv + 528 pp. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:SC

[Bri96v] Per Brinch Hansen. Shared classes. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 6, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:SOSa

[Bri96w] Per Brinch Hansen. Solo operating system: a Concurrent Pascal program. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 8, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:SOSb

[Bri96x] Per Brinch Hansen. Solo operating system: Processes, monitors, and classes. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 9, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:SM

[Bri96y] Per Brinch Hansen. Structured multiprogramming. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 4, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:SPL

[Bri96z] Per Brinch Hansen. SuperPascal: a publication language for parallel scientific computing. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 24, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1996:TMS

[Bri96-27] Per Brinch Hansen. Testing a multiprogramming system. In The Search for Simplicity: Essays in Parallel Programming [Bri96u], chapter 5, page ?? ISBN 0-8186-7566-7. LCCN QA76.642 .B756 1996.

BrinchHansen:1997:RSS

[Bri97] Per Brinch Hansen. Review of the search for simplicity essays in parallel programming. ACM SIGPLAN Notices, 32(11):31–??, November 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

BrinchHansen:1998:CGS

[Bri98a] Per Brinch Hansen. Conjugate gradient solution of linear equations. Concurrency: Practice and Experience, 10(2):139–156, February 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-

9128 (electronic). URL http://www3.interscience.wiley.com/
cgi-bin/abstract?ID=5370;   http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=5370&PLACEBO=IE.pdf.

BrinchHansen:1998:EHP

[Bri98b] Per Brinch Hansen. An evaluation of High Performance Fortran. ACM SIGPLAN Notices, 33(3):57–64, March 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). The author strongly criticizes HPF, and remarks “HPF has severe conceptual and practical limitations.” See response [Off98].

BrinchHansen:1998:EMP

[Bri98c] Per Brinch Hansen. An evaluation of the Message-Passing Interface. ACM SIGPLAN Notices, 33(3):65–72, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). The author criticizes MPI, and remarks “MPI … lack[s] the elegance and security that can only by checked by a parallel programming language.”.

BrinchHansen:1998:LE

[Bri98d] Per Brinch Hansen. Letter to the editor. ACM SIGPLAN Notices, 33(8):32–33, August 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

BrinchHansen:1999:C

[Bri99a] Per Brinch Hansen. Classes. In Programming for Everyone in Java [Bri99f], pages 183–221. ISBN 0-387-98683-9 (softcover), 1-4612-1514-5 (e-book). LCCN QA76.73.J38 B74 1999. US$34.95.

BrinchHansen:1999:JIP

[Bri99b] Per Brinch Hansen. Java’s insecure parallelism. ACM SIGPLAN Notices, 34(4):38–45, April 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

BrinchHansen:1999:Ma

[Bri99c] Per Brinch Hansen. Matrices. In Programming for Everyone in Java [Bri99f], pages 129–142. ISBN 0-387-98683-9 (softcover), 1-4612-1514-5 (e-book). LCCN QA76.73.J38 B74 1999. US$34.95.

BrinchHansen:1999:Mb

[Bri99d] Per Brinch Hansen. Methods. In Programming for Everyone in Java [Bri99f], pages 143–166. ISBN 0-387-98683-9 (softcover), 1-4612-1514-5 (e-book). LCCN QA76.73.J38 B74 1999. US$34.95.

BrinchHansen:1999:MC

[Bri99e] Per Brinch Hansen. Multiple choice. In Programming for Everyone in Java [Bri99f], pages 47–64. ISBN 0-387-98683-9 (softcover), 1-4612-1514-5 (e-book). LCCN QA76.73.J38 B74 1999. US$34.95.

BrinchHansen:1999:PEJ

[Bri99f] Per Brinch Hansen. Programming for Everyone in Java. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98683-9 (softcover), 1-4612-1514-5 (e-book). xviii + 238 pp. LCCN QA76.73.J38 B74 1999. US$34.95.

BrinchHansen:1999:R

[Bri99g] Per Brinch Hansen. Recursion. In Programming for Everyone in Java [Bri99f], pages 167–182. ISBN 0-387-98683-9 (softcover), 1-4612-1514-5 (e-book). LCCN QA76.73.J38 B74 1999. US$34.95.

BrinchHansen:1999:SA

[Bri99h] Per Brinch Hansen. Simple arrays. In Programming for Everyone in Java [Bri99f], pages 113–128. ISBN 0-387-98683-9 (softcover), 1-4612-1514-5 (e-book). LCCN QA76.73.J38 B74 1999. US$34.95.

BrinchHansen:1999:SC

[Bri99i] Per Brinch Hansen. Simple choice. In Programming for Everyone in Java [Bri99f], pages 27–45. ISBN 0-387-98683-9 (softcover), 1-4612-1514-5 (e-book). LCCN QA76.73.J38 B74 1999. US$34.95.

BrinchHansen:1999:SP

[Bri99j] Per Brinch Hansen. Simple programs. In Programming for Everyone in Java [Bri99f], pages 1–26. ISBN 0-387-98683-9 (softcover), 1-4612-1514-5 (e-book). LCCN QA76.73.J38 B74 1999. US$34.95.

BrinchHansen:1999:SR

[Bri99k] Per Brinch Hansen. Simple repetition. In Programming for Everyone in Java [Bri99f], pages 65–76. ISBN 0-387-98683-9 (softcover), 1-4612-1514-5 (e-book). LCCN QA76.73.J38 B74 1999. US$34.95.

BrinchHansen:1999:TCJ

[Bri99l] Per Brinch Hansen. Technical correspondence: Java’s insecure parallelism. ACM SIGPLAN Notices, 34(4):38–45, April 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print),

1558-1160 (electronic).  URL http://brinch-hansen.net/
papers/1999b.pdf;   http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/h/Hansen:Per_Brinch.html.

BrinchHansen:1999:TF

[Bri99m] Per Brinch Hansen. Text files. In Programming for Everyone in Java [Bri99f], pages 91–112. ISBN 0-387-98683-9 (softcover), 1-4612-1514-5 (e-book). LCCN QA76.73.J38 B74 1999. US$34.95.

BrinchHansen:1999:UR

[Bri99n] Per Brinch Hansen. Unbounded repetition. In Programming for Everyone in Java [Bri99f], pages 77–90. ISBN 0-387-98683-9 (softcover), 1-4612-1514-5 (e-book). LCCN QA76.73.J38 B74 1999. US$34.95.

BrinchHansen:2000:LEP

[Bri00] Per Brinch Hansen. Letter to the Editor: ‘Parallelism and recursion in message passing libraries: An efficient methodology’, C. Rodríguez, F. de Sande, C. León and L. García, Concurrency: Practice and Experience 1999; 11(7):355–365. Concurrency: Practice and Experience, 12(15):1511–1513, December 25, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-bin/
abstract/76505649/START;   http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=76505649&PLACEBO=IE.pdf.  See

[RDLG99, Leó00].

BrinchHansen:2001:COS

[Bri01a] Per Brinch Hansen. Classic Operating Systems: from Batch Processing to Distributed Systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 0-387-95113-X. x + 597 pp. LCCN QA76.76.O63

B7425 2001. URL http://www.loc.gov/catdir/enhancements/
fy0816/00045036-d.html;   http://www.loc.gov/catdir/enhancements/
fy0816/00045036-t.html;   https://link.springer.com/book/
10.1007/978-1-4757-3510-9.

BrinchHansen:2001:EOS

[Bri01b] Per Brinch Hansen. The evolution of operating systems. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 1–34. ISBN 0-387-95113-X. LCCN QA76.76.O63

B7425 2001. URL http://brinch-hansen.net/papers/2001b.
pdf.

BrinchHansen:2001:RSM

[Bri01c] Per Brinch Hansen. RC 4000 software: Multiprogramming system. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 237–281. ISBN 0-387-95113-X. LCCN QA76.76.O63 B7425 2001. URL http://www.loc.gov/

catdir/enhancements/fy0816/00045036-d.html;   http://www.
loc.gov/catdir/enhancements/fy0816/00045036-t.html;   https:
//link.springer.com/book/10.1007/978-1-4757-3510-9.

BrinchHansen:2001:SOSa

[Bri01d] Per Brinch Hansen. The Solo operating system: A Concurrent Pascal program. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 324–336. ISBN 0-387-95113-X. LCCN QA76.76.O63 B7425 2001. URL http://www.loc.gov/

catdir/enhancements/fy0816/00045036-d.html;   http://www.
loc.gov/catdir/enhancements/fy0816/00045036-t.html;   https:
//link.springer.com/book/10.1007/978-1-4757-3510-9.

BrinchHansen:2001:SOSb

[Bri01e] Per Brinch Hansen. The Solo operating system: Processes, monitors and classes. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 337–384. ISBN 0-387-95113-X. LCCN QA76.76.O63 B7425 2001. URL http://www.

loc.gov/catdir/enhancements/fy0816/00045036-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0816/00045036-t.html;
 https://link.springer.com/book/10.1007/978-1-4757-3510-
9.

BrinchHansen:2002:DP

[Bri02a] Per Brinch Hansen. Design principles. In The Origin of Concurrent Programming: From Semaphores to Remote Procedure Calls [Bri02f], pages 382–393. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN QA76.642 .O75 2002. URL

http://link.springer.com/10.1007/978-1-4757-3472-0.

BrinchHansen:2002:DPC

[Bri02b] Per Brinch Hansen. Distributed processes: a concurrent programming concept. In The Origin of Concurrent Programming: From Semaphores to Remote Procedure Calls [Bri02f], pages 444–463. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7.

LCCN QA76.642 .O75 2002. URL http://link.springer.com/
10.1007/978-1-4757-3472-0.

BrinchHansen:2002:EPR

[Bri02c] Per Brinch Hansen. Efficient parallel recursion. In The Origin of Concurrent Programming: From Semaphores to Remote Procedure Calls [Bri02f], pages 525–534. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN QA76.642 .O75 2002. URL

http://link.springer.com/10.1007/978-1-4757-3472-0.

BrinchHansen:2002:ICP

[Bri02d] Per Brinch Hansen. The invention of concurrent programming. In The Origin of Concurrent Programming: From Semaphores to Remote Procedure Calls [Bri02f], pages 3–61. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN QA76.642 .O75

2002. URL http://brinch-hansen.net/papers/2002b.pdf.

BrinchHansen:2002:JPL

[Bri02e] Per Brinch Hansen. Joyce — a programming language for distributed systems. In The Origin of Concurrent Programming: From Semaphores to Remote Procedure Calls [Bri02f], pages 464–492. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN QA76.642 .O75 2002. URL http://link.

springer.com/10.1007/978-1-4757-3472-0.

BrinchHansen:2002:OCP

[Bri02f] Per Brinch Hansen, editor. The Origin of Concurrent Programming: From Semaphores to Remote Procedure Calls. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. x + 534 pp. LCCN QA76.642 .O75 2002. URL

http://link.springer.com/10.1007/978-1-4757-3472-0.

BrinchHansen:2002:OCO

[Bri02g] Per Brinch Hansen. An outline of a course on operating system principles. In The Origin of Concurrent Programming: From Semaphores to Remote Procedure Calls [Bri02f], pages 245–254. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7.

LCCN QA76.642 .O75 2002. URL http://link.springer.com/
10.1007/978-1-4757-3472-0.

BrinchHansen:2002:PLC

[Bri02h] Per Brinch Hansen. The programming language Concurrent Pascal. In The Origin of Concurrent Programming: From Semaphores to Remote Procedure Calls [Bri02f], pages 297–318. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN

QA76.642 .O75 2002. URL http://link.springer.com/10.
1007/978-1-4757-3472-0.

BrinchHansen:2002:RSM

[Bri02i] Per Brinch Hansen. RC 4000 software: Multiprogramming system. In The Origin of Concurrent Programming: From Semaphores to Remote Procedure Calls [Bri02f], pages 153–197. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN

QA76.642 .O75 2002. URL http://link.springer.com/10.
1007/978-1-4757-3472-0.

BrinchHansen:2002:SC

[Bri02j] Per Brinch Hansen. Shared classes. In The Origin of Concurrent Programming: From Semaphores to Remote Procedure Calls [Bri02f], pages 265–271. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN QA76.642 .O75 2002. URL

http://link.springer.com/10.1007/978-1-4757-3472-0.

BrinchHansen:2002:SOSa

[Bri02k] Per Brinch Hansen. The Solo operating system: a Concurrent Pascal program. In The Origin of Concurrent Programming: From Semaphores to Remote Procedure Calls [Bri02f], pages 321–333. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7.

LCCN QA76.642 .O75 2002. URL http://link.springer.com/
10.1007/978-1-4757-3472-0.

BrinchHansen:2002:SOSb

[Bri02l] Per Brinch Hansen. The Solo operating system: Processes, monitors and classes. In The Origin of Concurrent Programming: From Semaphores to Remote Procedure Calls [Bri02f], pages 334–381. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7.

LCCN QA76.642 .O75 2002. URL http://link.springer.com/
10.1007/978-1-4757-3472-0.

BrinchHansen:2002:SM

[Bri02m] Per Brinch Hansen. Structured multiprogramming. In The Origin of Concurrent Programming: From Semaphores to Remote Procedure Calls [Bri02f], pages 255–264. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN QA76.642 .O75

2002. URL http://link.springer.com/10.1007/978-1-4757-
3472-0.

BrinchHansen:2002:SPL

[Bri02n] Per Brinch Hansen. SuperPascal: a publication language for parallel scientific computing. In The Origin of Concurrent Programming: From Semaphores to Remote Procedure Calls [Bri02f], pages 495–524. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN QA76.642 .O75 2002. URL http://link.

springer.com/10.1007/978-1-4757-3472-0.

BrinchHansen:2004:PS

[Bri04] Per Brinch Hansen. A programmer’s story. Web site., 2004. URL

http://brinch-hansen.net/.

Bauer:1976:LHI

[BS76] Friedrich Ludwig Bauer and Klaus Samelson, editors. Language hierarchies and interfaces: international summer school [held July 23 to August 2, 1975 in Marktoberdorf], volume 46 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. ISBN 0-387-07994-7 (New York), 3-540-07994-7 (Berlin), 3-540-37972-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6

.L335. URL http://link.springer-ny.com/link/service/
series/0558/tocs/t0046.htm;   http://www.springerlink.
com/content/978-0-387-07994-3;   http://www.springerlink.
com/openurl.asp?genre=book&isbn=978-3-540-07994-1;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=0302-
9743&volume=46.

BrinchHansen:1978:SIM

[BS78] Per Brinch Hansen and Jørgen Staunstrup. Specification and implementation of mutual exclusion. IEEE Transactions on Software Engineering, SE-4(5):365–370, September 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Courtois:1972:CCT

[CHP72] P. J. Courtois, F. Heymans, and David Lorge Parnas. Comments on A Comparison of Two Synchronizing Concepts by Per Brinch Hansen. Acta Informatica, 1(??):375–376, ???? 1972. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [Bri72a, Bri73b].

Corbato:2001:ETS

[CMDD01] Fernando J. Corbató, Marjorie Merwin-Daggett, and Robert C. Daley. An experimental time-sharing system. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 117–137. ISBN 0-387-95113-X. LCCN

QA76.76.O63 B7425 2001. URL http://www.loc.gov/catdir/
enhancements/fy0816/00045036-d.html;   http://www.loc.
gov/catdir/enhancements/fy0816/00045036-t.html;   https:
//link.springer.com/book/10.1007/978-1-4757-3510-9.

Dijkstra:1989:LSR

[DF89] Edsger W. Dijkstra and W. H. J. Feijen. The linear search revisited. Structured Programming, 10(1):5–9, 1989. CODEN STPGEM. ISSN 0935-1183.

Dijkstra:2001:SMS

[Dij01] Edsger W. Dijkstra. The structure of the “THE” multiprogramming system. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 223–236. ISBN 0-387-95113-X. LCCN QA76.76.O63 B7425 2001. URL http://www.loc.gov/

catdir/enhancements/fy0816/00045036-d.html;   http://www.
loc.gov/catdir/enhancements/fy0816/00045036-t.html;   https:
//link.springer.com/book/10.1007/978-1-4757-3510-9.

Dijkstra:2002:CSP

[Dij02a] Edsger W. Dijkstra. Cooperating sequential processes. In Brinch Hansen [Bri02f], pages 65–138. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN QA76.642 .O75

2002. URL http://link.springer.com/10.1007/978-1-4757-
3472-0.

Dijkstra:2002:HOS

[Dij02b] Edsger W. Dijkstra. Hierarchical ordering of sequential processes. In Brinch Hansen [Bri02f], pages 198–227. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN QA76.642 .O75

2002. URL http://link.springer.com/10.1007/978-1-4757-
3472-0.

Dijkstra:2002:SMS

[Dij02c] Edsger W. Dijkstra. The structure of the “THE” multiprogramming system. In Brinch Hansen [Bri02f], pages 139–152. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7.

LCCN QA76.642 .O75 2002. URL http://link.springer.com/
10.1007/978-1-4757-3472-0.

Dijkstra:2002:SE

[Dij02d] Edsger W. Dijkstra. A synthesis emerging? In Brinch Hansen [Bri02f], pages 397–412. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN QA76.642 .O75 2002. URL

http://link.springer.com/10.1007/978-1-4757-3472-0.

Daley:2001:GPF

[DN01] Robert C. Daley and Peter G. Neumann. A general-purpose file system for secondary storage. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 138–166. ISBN 0-387-95113-X. LCCN QA76.76.O63

B7425 2001. URL http://www.loc.gov/catdir/enhancements/
fy0816/00045036-d.html;   http://www.loc.gov/catdir/enhancements/
fy0816/00045036-t.html;   https://link.springer.com/book/
10.1007/978-1-4757-3510-9.

Freiman:1971:IPP

[FGR71] Charles V. Freiman, John E. Griffith, and Jack L. Rosenfeld, editors. Information Processing, Proceedings of IFIP Congress 1971, Volume 1 — Foundations and Systems, Ljubljana, Yugoslavia, August 23–28, 1971. North-Holland Publishing Co., Amsterdam, The Netherlands, 1971. ISBN 0-7204-2063-6 (two volumes). LCCN QA75.5 .I532 1971, TK7885.A1 I61 1971.

Fraser:2001:FID

[Fra01] A. G. Fraser. File integrity in a disc-based multi-access system. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 167–194. ISBN 0-387-95113-X. LCCN

QA76.76.O63 B7425 2001. URL http://www.loc.gov/catdir/
enhancements/fy0816/00045036-d.html;   http://www.loc.
gov/catdir/enhancements/fy0816/00045036-t.html;   https:
//link.springer.com/book/10.1007/978-1-4757-3510-9.

Hoare:2002:CSP

[Hoa02a] C. A. R. Hoare. Communicating sequential processes. In Brinch Hansen [Bri02f], pages 413–443. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN QA76.642 .O75

2002. URL http://link.springer.com/10.1007/978-1-4757-
3472-0.

Hoare:2002:MOS

[Hoa02b] C. A. R. Hoare. Monitors: An operating system structuring concept. In Brinch Hansen [Bri02f], pages 272–294. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN QA76.642

.O75 2002. URL http://link.springer.com/10.1007/978-1-
4757-3472-0.

Hoare:2002:TTP

[Hoa02c] C. A. R. Hoare. Towards a theory of parallel programming. In Brinch Hansen [Bri02f], pages 231–244. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN QA76.642 .O75

2002. URL http://link.springer.com/10.1007/978-1-4757-
3472-0.

Horspool:2007:OPB

[Hor07] Nigel Horspool. Obituary: Per Brinch Hansen (1938–2007). Software—Practice and Experience, 37(12):1243, October 2007. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kilburn:2001:AS

[KPH01] Tom Kilburn, R. Bruce Payne, and David J. Howarth. The Atlas supervisor. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 49–77. ISBN 0-387-95113-X. LCCN QA76.76.O63 B7425 2001. URL http://www.loc.gov/

catdir/enhancements/fy0816/00045036-d.html;   http://www.
loc.gov/catdir/enhancements/fy0816/00045036-t.html;   https:
//link.springer.com/book/10.1007/978-1-4757-3510-9.

Kruijer:1997:BNS

[Kru97] H. S. M. Kruijer. Book nook: The Search for Simplicity Essays in Parallel Programming by Per Brinch Hansen. ACM SIGPLAN Notices, 32(11):31–32, November 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lauesen:2001:LSB

[Lau01] Søren Lauesen. A large semaphore based operating system. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 295–323. ISBN 0-387-95113-X. LCCN

QA76.76.O63 B7425 2001. URL http://www.loc.gov/catdir/
enhancements/fy0816/00045036-d.html;   http://www.loc.
gov/catdir/enhancements/fy0816/00045036-t.html;   https:
//link.springer.com/book/10.1007/978-1-4757-3510-9.

Leon:2000:ARP

[Leó00] Casiano Rodríguez León. Authors’ response: ‘Parallelism and recursion in message passing libraries: An efficient methodology’, C. Rodríguez, F. de Sande, C. León and L. García, Concurrency: Practice and Experience 1999; 11(7):355–365. Concurrency: Practice and Experience, 12(15):1515–1516, December 25, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-bin/
abstract/76505647/START;   http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=76505647&PLACEBO=IE.pdf.  See

[RDLG99, Bri00].

Lee:2003:BPB

[LH03] J. A. N. Lee and Thomas Haigh. Biographies: Per Brinch Hansen. IEEE Annals of the History of Computing, 25(1):80–83, January/ March 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Liskov:2001:DVO

[Lis01] Barbara H. Liskov. The design of the Venus operating system. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 282–294. ISBN 0-387-95113-X. LCCN

QA76.76.O63 B7425 2001. URL http://www.loc.gov/catdir/
enhancements/fy0816/00045036-d.html;   http://www.loc.
gov/catdir/enhancements/fy0816/00045036-t.html;   https:
//link.springer.com/book/10.1007/978-1-4757-3510-9.

Linger:1979:SPT

[LMW79] R. C. Linger, H. D. Mills, and B. I. Witt. Structured Programming: Theory and Practice. Addison-Wesley, Reading, MA, USA, 1979. ISBN 0-201-14461-1. xi + 402 pp. LCCN QA76.6 .L55.

Lampson:2001:OOS

[LS01] Butler W. Lampson and Robert F. Sproull. An open operating system for a single-user machine. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 414–432. ISBN 0-387-95113-X. LCCN

QA76.76.O63 B7425 2001. URL http://www.loc.gov/catdir/
enhancements/fy0816/00045036-d.html;   http://www.loc.
gov/catdir/enhancements/fy0816/00045036-t.html;   https:
//link.springer.com/book/10.1007/978-1-4757-3510-9.

Lynch:2001:DHC

[Lyn01] William C. Lynch. Description of a high capacity, fast turnaround university computing center. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 88–101. ISBN 0-387-95113-X. LCCN

QA76.76.O63 B7425 2001. URL http://www.loc.gov/catdir/
enhancements/fy0816/00045036-d.html;   http://www.loc.

gov/catdir/enhancements/fy0816/00045036-t.html;   https:
//link.springer.com/book/10.1007/978-1-4757-3510-9.

Naur:1975:NBJ

[Nau75] P. Naur. New books and journals. BIT, 15(4):455–457, December 1975. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print),

1572-9125 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=0006-3835&volume=15&issue=

4&spage=455. Includes review of Operating Systems Principles, by Per Brinch Hansen.

Offner:1998:PBH

[Off98] Carl D. Offner. Per Brinch Hansen’s concerns about High Performance Fortran. ACM SIGPLAN Notices, 33(8):34–39, August 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [Bri98b].

Oliphint:2001:OSB

[Oli01] Clark Oliphint. Operating system for the B 5000. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 78–87. ISBN 0-387-95113-X. LCCN

QA76.76.O63 B7425 2001. URL http://www.loc.gov/catdir/
enhancements/fy0816/00045036-d.html;   http://www.loc.
gov/catdir/enhancements/fy0816/00045036-t.html;   https:
//link.springer.com/book/10.1007/978-1-4757-3510-9.

Piner:2002:CSCf

[Pin02] Mary-Louise G. Piner. Computer Society connection: IEEE Computer Society recognizes achievements of leading computer engineers: Per Brinch Hansen recognized for pioneering works; Stanford President Hennessy receives architecture award; Ian F. Akyildiz named Goode winner; Society digital library collection is expanded; Mills Award honors software engineering achievements. Computer, 35(6):74–76, June 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/dl/mags/co/2002/06/r6074.htm;
 http://csdl.computer.org/dl/mags/co/2002/06/r6074.pdf.

Redell:2001:POS

[RDH^{+}01] David D. Redell, Yogen K. Dalal, Thomas R. Horsley, Hugh C. Lauer, William C. Lynch, Paul R. McJones, Hal G. Murray, and Stephen C. Purcell. Pilot: an operating system for a personal computer. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 433–459. ISBN 0-387-95113-X. LCCN QA76.76.O63

B7425 2001. URL http://www.loc.gov/catdir/enhancements/
fy0816/00045036-d.html;   http://www.loc.gov/catdir/enhancements/
fy0816/00045036-t.html;   https://link.springer.com/book/
10.1007/978-1-4757-3510-9.

Rodriguez:1999:PRM

[RDLG99] C. Rodríguez, F. De Sande, C. León, and L. García. Parallelism and recursion in message passing libraries: an efficient methodology. Concurrency: Practice and Experience, 11(7):355– 365, June 1999. CODEN CPEXEI. ISSN 1040-3108 (print),

1096-9128 (electronic).  URL http://www3.interscience.
wiley.com/cgi-bin/abstract?ID=62500298;   http://www3.
interscience.wiley.com/cgi-bin/fulltext?ID=62500298&PLACEBO=

IE.pdf. See correspondence [Bri00, Leó00].

Regener:1984:MID

[Reg84] Eric Regener. Multiprecision integer division examples using arbitrary radix. ACM Transactions on Mathematical Software, 10(3): 325–328, September 1984. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [Bri94c].

Ritchie:2001:UTS

[RT01] Dennis M. Ritchie and Ken Thompson. The Unix time-sharing system. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 195–220. ISBN 0-387-95113-X. LCCN QA76.76.O63 B7425 2001. URL http://www.loc.gov/

catdir/enhancements/fy0816/00045036-d.html;   http://www.
loc.gov/catdir/enhancements/fy0816/00045036-t.html;   https:
//link.springer.com/book/10.1007/978-1-4757-3510-9.

Ryckman:2001:ICG

[Ryc01] George F. Ryckman. The IBM 701 computer at the General Motors Research Laboratories. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 37–40. ISBN 0-387-95113-X. LCCN QA76.76.O63

B7425 2001. URL http://www.loc.gov/catdir/enhancements/
fy0816/00045036-d.html;   http://www.loc.gov/catdir/enhancements/
fy0816/00045036-t.html;   https://link.springer.com/book/
10.1007/978-1-4757-3510-9.

Smith:2001:SUI

[SIKH01] David C. Smith, Charles Irby, Ralph Kimball, and Eric Harslem. The Star user interface: an overview. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 460–490. ISBN 0-387-95113-X. LCCN

QA76.76.O63 B7425 2001. URL http://www.loc.gov/catdir/
enhancements/fy0816/00045036-d.html;   http://www.loc.
gov/catdir/enhancements/fy0816/00045036-t.html;   https:
//link.springer.com/book/10.1007/978-1-4757-3510-9.

Swinehart:2001:WSS

[SMB01] Daniel Swinehart, Gene McDaniel, and David Boggs. WFS: a simple shared file system for a distributed environment. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 493–510. ISBN 0-387-95113-X. LCCN

QA76.76.O63 B7425 2001. URL http://www.loc.gov/catdir/
enhancements/fy0816/00045036-d.html;   http://www.loc.
gov/catdir/enhancements/fy0816/00045036-t.html;   https:
//link.springer.com/book/10.1007/978-1-4757-3510-9.

Smith:2001:BSP

[Smi01] Richard B. Smith. The BKS system for the Philco-2000. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 43–45. ISBN 0-387-95113-X. LCCN

QA76.76.O63 B7425 2001. URL http://www.loc.gov/catdir/
enhancements/fy0816/00045036-d.html;   http://www.loc.
gov/catdir/enhancements/fy0816/00045036-t.html;   https:
//link.springer.com/book/10.1007/978-1-4757-3510-9.

Shrivastava:2001:DRR

[SP01] Santosh Shrivastava and Fabio Panzieri. The design of a reliable remote procedure call mechanism. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 511–527. ISBN 0-387-95113-X. LCCN

QA76.76.O63 B7425 2001. URL http://www.loc.gov/catdir/
enhancements/fy0816/00045036-d.html;   http://www.loc.
gov/catdir/enhancements/fy0816/00045036-t.html;   https:
//link.springer.com/book/10.1007/978-1-4757-3510-9.

Stoy:2001:EOS

[SS01] Joe E. Stoy and Christopher Strachey. OS 6 — an experimental operating system for a small computer: Input/output and filing system. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 387–413. ISBN 0-387-95113-X. LCCN QA76.76.O63 B7425 2001. URL http://www.loc.gov/

catdir/enhancements/fy0816/00045036-d.html;   http://www.
loc.gov/catdir/enhancements/fy0816/00045036-t.html;   https:
//link.springer.com/book/10.1007/978-1-4757-3510-9.

Tanenbaum:2001:EAD

[TRS^{+}01] Andrew S. Tanenbaum, Robbert Van Renesse, Hans Van Staveren, Gregory J. Sharp, Sape J. Mullender, Jack Jansen, and Guido Van Rossum. Experiences with the Amoeba distributed operating system. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01a], pages 550–586. ISBN 0-387-95113-X. LCCN QA76.76.O63 B7425 2001. URL http://www.loc.gov/

catdir/enhancements/fy0816/00045036-d.html;   http://www.
loc.gov/catdir/enhancements/fy0816/00045036-t.html;   https:
//link.springer.com/book/10.1007/978-1-4757-3510-9.

VandeVelde:1995:BNR

[Van95] Erik F. Van de Velde. Book news & reviews: Studies in Computational Science: Parallel Programming Paradigms by Per Brinch Hansen. IEEE Computational Science & Engineering, 2(4):85–87, Winter 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Varma:2002:BSB

[VBM^{+}02] Tathagat Varma, Fernando Berzal, Nicolás Marín, Tope Omitola, and Matt Edgar. Bookshelf: Strategies for building better people; no bulletproof vest for your system; learning from the past; a treasure for software designers. IEEE Software, 19(6):101–104, November/December 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://dlib.computer.

org/so/books/so2002/pdf/s6101.pdf. Review of The People-

ware Papers: Notes on the Human Side of Software by Larry L. Constantine, Fundamentals of Network Security by John E. Canavan, Classic Operating Systems by Per Brinch Hansen, and Software Fundamentals: Collected Papers by David L. Parnas by David M. Weiss and Daniel M. Hoffman, eds.

Whiteside:1978:BRR

[Whi78] A. E. Whiteside. Book reviews: Review of The Architecture of Concurrent Programs by Per Brinch Hansen, Prentice-Hall 1977. ACM SIGARCH Computer Architecture News, 6(6):32, February REFERENCES 44

1978. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).